OpenStudy (anonymous):

I will medal and fan you!!! Please help How do I figure this out? Tom has eight red balls. He has two less balls then green balls. He has 189 more green balls then red balls. He has 12 pink marbles. How many balls does he have ALTOGETHER?

OpenStudy (anonymous):

This is so confusing.

OpenStudy (anonymous):

I know :/

OpenStudy (anonymous):

Pink: 12 Green 189 more than red, but red had 8..so whutt

OpenStudy (anonymous):

Its 217

OpenStudy (anonymous):

That's like really really easy

OpenStudy (anonymous):

*he has two blue less then green Sorry! I was typing, it wouldn't let me copy it

OpenStudy (anonymous):

@KenichiRender Please explain why :)

OpenStudy (anonymous):

2 bkue less then green, okay hold up

OpenStudy (mathmath333):

r=8 b=g-2 g=r+189 p=12

OpenStudy (anonymous):

its 402

OpenStudy (anonymous):

total

OpenStudy (anonymous):

197 + 195 + 12 + 8 = 402

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

mhm